OpenStudy (anonymous):

Which one is a rule for the nth term of an arithmetic sequence with a8 = 56 and d = 5?

OpenStudy (anonymous):

\[a _{8}=56 \] and d=5

OpenStudy (anonymous):

Select one of the options below as your answer: A. an = 16 − 5n B. an = 26 + 5n C. an = 26 − 5n D. an = 16 + 5n

OpenStudy (anonymous):

D

OpenStudy (akshay_budhkar):

how?

OpenStudy (akshay_budhkar):

an=a+(n-1)d

OpenStudy (akshay_budhkar):

therefore a8=a+7d

OpenStudy (akshay_budhkar):

56=a+35

OpenStudy (akshay_budhkar):

a=21

OpenStudy (akshay_budhkar):

an=21+(n-1)5

OpenStudy (akshay_budhkar):

an=16+5n

OpenStudy (akshay_budhkar):

so d ishaan is right:)